# PACE: Physics Informed Uncertainty Aware Climate Emulator

### v1

@@ Abstract
Climate models serve as critical tools for evaluating the effects of climate change and projecting future climate scenarios. However, the reliance on numerical simulations of physical equations renders them computationally intensive and inefficient. While deep learning methodologies have made significant progress in weather forecasting, they are still unstable for climate emulation tasks. Here, we propose PACE, a lightweight 684K parameter Physics Informed Uncertainty Aware Climate Emulator. PACE emulates temperature and precipitation stably for 86 years while only being trained on greenhouse gas emissions data. We incorporate a fundamental physical law of advection-diffusion in PACE accounting for boundary conditions and empirically estimating the diffusion co-efficient and flow velocities from emissions data. PACE has been trained on 15 climate models provided by ClimateSet outperforming baselines across most of the climate models and advancing a new state of the art in a climate diagnostic task.
@@ 1 Introduction
To address this gap, we propose PACE, which treats climate emulation as a diagnostic-type prediction and integrate emissions data directly into the model’s training framework, to predict climate variables from a given parallel time-series of climate forcer emission maps (GHG and aerosols) allowing for more accurate simulation of future climate states under varying concentration scenarios.
Furthermore, we focus on two key phenomenon observed by our climate system i.e. advection and diffusion. In climate modeling, the advection-diffusion equation is fundamental for simulating the transport and dispersion of climate variables, such as temperature and moisture (Choi et al., 2023). PACE proves to be compute-efficient by modelling key physical law which reduces its dependence on large datasets making it data-efficient 1. Additionally, with generalizability inherent in the advection-diffusion equation, PACE generalizes across 15 climate models emulating surface temperature and precipitation stably for 86 years solely from emissions data (see Figure 1). Our contributions are as follows:
@@
    We propose PACE, a Neural ODE based climate emulator which models the advection-diffusion phenomenon by dynamically estimating the diffusion coefficient and flow velocities based on the input greenhouse gas concentrations.
@@
    Finally, we perform extensive experiments to show the generalization capabilities of PACE for emulating 15 climate models for 86 years at one time.
@@
Table 1: Computational Efficiency of PACE vs several Climate Emulators
@@ 3.3 Convolution Block Attention Module (CBAM)
Figure 2: Complete architectural pipeline of PACE. The model initializes the velocity fields and diffusion co-efficient from the input data of four gases i.e. $C0_{2},CH_{4},BC,SO_{2}$. After solving the advection-diffusion equation, channel and spatial attention extracts important spatial features to generate output maps of Temperature and Precipitation
@@ 4.1 Task
The goal of PACE is to emulate surface air temperature (TAS) and precipitation (PR) from climate forcer emission maps (CO2, CH4, SO2, BC) for a parallel time series of 2015-2100. We simulate the output of each climate model as single and super emulator, and also validate the generalisation of our methodology using zero-shot learning. We compare PACE against all baselines provided by ClimateSet under the same hyperparameter settings. We also compare against ACE (Watt-Meyer et al., 2023) and LUCIE Guan et al. (2024), two recent climate emulators. Since, they both are developed for different emulation task, we adopt their base architecture SFNO Bonev et al. (2023) and train it for the same task as ours. The details for adaptation of SFNO are given in Appendix A.2.
@@ 4.2 Dataset
We train PACE on a total of 15 climate models provided by ClimateSet Kaltenborn et al. (2023). ClimateSet compiles climate data from the Coupled Model Intercomparison Project Phase 6 (CMIP6) (Eyring et al., 2016) , incorporating climate model outputs from ScenarioMIP (O’Neill et al., 2016) and future emission trajectories of climate forcing agents from Input Datasets for Model Intercomparison Projects (Input4MIPs) (Durack et al., 2017). Each climate model has been standardized to a spatial resolution of 250km i.e. 96 $\times$ 144 grid points (latitude $\times$ longitude) with a monthly temporal resolution. Both input and output datasets consist of 86-year time-series data spanning four SSP scenarios (SSP1-2.6, SSP2-4.5, SSP3-7.0, SSP5-8.5) from 2015 to 2100. We use three scenarios namely SSP1-2.6, SSP3-7.0, SSP5-8.5 for training with a validation split of 0.1 and SSP2-4.5 for testing.
@@ 4.3 Evaluation Metrics
We evaluate PACE and all benchmarks using latitude-weighted Root Mean Square Error (RMSE) given in equation 19.
@@ 4.4 Single Emulator
For single emulator experiments, we trained all models for 25 epochs. We report RMSE for UNet, ConvLSTM, ClimaX, ClimaX_Frozen and SFNO. The training hyperparameters are all kept similar to those used in the original paper. PACE outperforms all models for emulating temperature across 13 climate models whereas ClimaX performs better on EC-Earth3 and TaiESM1 with lowest RMSE. For precipitation emulation, all models perform slightly worse with SFNO having lowest RMSE for simulating BCC-CSM2-MR and TaiESM1, ClimaX for AWI-CM-1-1-MR and CAS-ESM2-0, UNet performs best for INM-CM4-8 while PACE performs best for simulating remaining 10 climate models. Figure 3 shows the overall distribution of how each ML model performs across all 15 climate models in emulating temperature and precipitation. Detailed RMSE results for all 15 climate models for temperature and precipitation emulation are given in Appendix B in Figures 6 and 7.
Figure 3: RMSE distribution for climate variable predictions ccross different models. Each box represents the interquartile range (IQR) of RMSE values, capturing the spread and variability in prediction accuracy for key climate variables Temperature and Precipitation. The plot shows the consistent performance of PACE across all 15 climate models.
@@ 4.5 Super Emulator
Here, the term super emulator is used to train a single ML model on all of 15 climate models. This leads to the rich feature learning resulting in better generalization capabilities across different climate models. We use the same multihead decoder proposed in ClimateSet to train all ML models including PACE. We train all ML models for 100 epochs to keep the training regime computationally efficient with 2 convolutional layers and and 32 units decoder head. For super-emulator experiments we use a batch size of 1 for all models due to computational constraints.
For super emulation PACE outperforms all ML models on 13 climate models while ConvLSTM performs best for emulating EC-Earth3-Veg-LR and TaiESM1. The authors of ClimateSet Kaltenborn et al. (2023) suggest that during super emulation, smaller models demonstrate superior learning efficiency compared to larger models. This is because smaller models converge faster, allowing the model to learn patterns and relationships in the data more rapidly. We believe PACE being physically consistent and compute-efficient is able to learn complex climate features and outperform computationally intensive climate emulators.
@@
Climate Models PACE UNet ConvLSTM SFNO ClimaX ClimaXfrozen
@@
4.6 Generalization Capabilities of PACE as a Single Emulator
We perform generalization experiments to test the generalization capabilities of ML models on three climate models: AWI-CM-1-1-MR, MPI-ESM1-2-HR and FGOALS-f3-L. The RMSE results are shown in Tables 3, 4 and 5 for TAS (surface air temperature) and PR (precipitation) pre-trained on different climate models and tested on these three climate models. The metric for best performing model is emboldened. The pretrained climate model column shows which dataset the model was initially trained on before being tested on the either of the three climate models. Overall we observe that PACE, SFNO and ClimaX performs well for generalizing over different climate models with PACE outperforming on majority of the models. Here ClimaX does gain benefit from being pre-trained on a range of climate models initially. In future, we will test ClimaX generalization without using the pre-trained checkpoint to see how it compares against the smaller models such as unet and convlstm.
@@ 4.6 Generalization Capabilities of PACE as a Single Emulator
Pre-Trained Climate Model PACE UNet ConvLSTM SFNO ClimaX ClimaXfrozen TAS PR TAS PR TAS PR TAS PR TAS PR TAS PR AWI-CM-1-1-MR 0.184 0.499 0.289 0.571 0.451 0.622 0.200 0.501 0.207 0.498 0.412 0.707 BCC-CSM2-MR 0.247 0.620 0.275 0.653 0.466 0.696 0.230 0.599 0.232 0.617 0.432 0.753 CAS-ESM2-0 0.275 0.641 0.273 0.694 0.477 0.714 0.263 0.656 0.250 0.654 0.461 0.763 MRI-ESM2-0 0.205 0.620 0.276 0.656 0.456 0.697 0.218 0.635 0.223 0.643 0.410 0.756 NorESM2-MM 0.231 0.582 0.301 0.562 0.459 0.674 0.245 0.591 0.286 0.551 0.441 0.754
@@
Pre-Trained Climate Model PACE UNet ConvLSTM SFNO ClimaX ClimaXfrozen TAS PR TAS PR TAS PR TAS PR TAS PR TAS PR MPI-ESM1-2-HR 0.193 0.477 0.234 0.410 0.449 0.636 0.198 0.586 0.214 0.509 0.411 0.716 CNRM-CM6-1-HR 0.221 0.580 0.268 0.657 0.482 0.714 0.234 0.601 0.225 0.599 0.453 0.756 EC-Earth3 0.228 0.544 0.262 0.558 0.465 0.667 0.259 0.551 0.220 0.567 0.443 0.744 EC-Earth3-Veg-LR 0.241 0.554 0.270 0.560 0.471 0.660 0.224 0.515 0.233 0.551 0.456 0.744 TaiESM1 0.208 0.602 0.294 0.711 0.461 0.692 0.248 0.645 0.268 0.682 0.427 0.757
@@
Pre-Trained Climate Model PACE UNet ConvLSTM SFNO ClimaX ClimaXfrozen TAS PR TAS PR TAS PR TAS PR TAS PR TAS PR FGOALS-f3-L 0.184 0.559 0.241 0.562 0.485 0.652 0.253 0.561 0.218 0.573 0.456 0.729 GFDL-ESM4 0.207 0.563 0.321 0.716 0.484 0.697 0.271 0.600 0.330 0.708 0.468 0.762 INM-CM4-8 0.232 0.722 0.296 0.776 0.491 0.744 0.259 0.737 0.245 0.745 0.468 0.790 INM-CM5-0 0.212 0.574 0.277 0.756 0.488 0.739 0.261 0.730 0.250 0.725 0.459 0.785 MPI-ESM1-2-HR 0.209 0.696 0.283 0.701 0.482 0.714 0.257 0.701 0.250 0.691 0.445 0.760
@@
We utilize FDM for spatial discretization in PACE which divides the the physical space (in this case Earth atmosphere) into a grid of discrete points. Each grid point represents a specific location, and the value of the physical quantity (e.g., temperature) is computed at each point. The gridding at a lower resolution does induce additional errors. In future, we aim to test FVM and FEM to test if they results in smoother outputs and reduce errors.
@@ 6 Ablation Studies
To understand the importance of each component of PACE, we perform ablation studies across four climate models namely AWI-CM-1-1-MR, TaiESM1, EC-Earth3 and NorESM2-MM.
Advection-Only: For this study, we remove the empirical estimated diffusion term from PACE and only model the advection process using Neural ODE. The resulting RMSE for surface air temperature and precipitation increases deteriorating the model’s overall performance. The results show that missing approximation of diffusion has a greater effect on temperature as compared to precipitation, therefore determining that diffusion is critical in accurately simulating the transport of physical quantities like heat, moisture, and momentum.
@@ 7 Conclusion and Future Work
In this work, we present PACE, a physics and uncertainty aware climate emulator which accounts for Earth’s atmospheric advection-diffusion phenomenon. We incorporate a key physical law in PACE by solving a time-dependent partial different equation (PDE) using Neural ODE. Additionally, we encode periodic boundary conditions to avoid artificial edge effects that arise from rigid boundaries.
However, there are a few potential limitations of PACE which will be addressed as a possible future work. Our training regime only includes climate model data which is based on simulations and not considered entirely accurate. As a future work, we aim to extend training mechanism of PACE on both ERA5 weather and ClimateSet’s extensive climate data to enhance emulation accuracy. Additionally, PACE is trained on coarse resolution data which does not fully account for extreme events at regional level. Further, PACE is still limited in its ability to emulate precipitation stably for 86 years. These limitations can be addressed by training on high resolution data and encoding physical constraints such as energy, mass and water conservation in a loss function.
@@ A.1 Hyperparameters
Table 6: Training Hyperparameters for PACE
@@ A.3 Hardware and Software Requirements
We use PyTorch Paszke et al. (2019), Pytorch Lightning Falcon (2019), torchdiffeq Chen et al. (2018) for implementation of PACE. We train PACE on a single RTXA5000 with 24GB RAM. We perform all super emulator training experiments on a single NVIDIA DGX A100 with 80 GB RAM.
@@ Appendix B Results
Figure 6: RMSE results for Surface Air Temperature (TAS) Emulation for the projection of SSP2-4.5 (2015 – 2100). PACE outperforms all other ML models on 13 out of 15 Climate models namely:
Figure 7: RMSE results for Precipitation (PR) Emulation for the projection of SSP2-4.5 (2015 – 2100). PACE outperforms all other ML models on 9 out of 15 Climate models namely:

### v2

# PACER: Physics Informed Uncertainty Aware Climate EmulatoR

@@ Abstract
Climate models serve as critical tools for evaluating the effects of climate change and projecting future climate scenarios. However, the reliance on numerical simulations of physical equations renders them computationally intensive and inefficient. While deep learning methodologies have made significant progress in weather forecasting, they are still unstable for climate emulation tasks. Here, we propose PACER, a lightweight 684K parameter Physics Informed Uncertainty Aware Climate EmulatoR. PACER emulates temperature and precipitation stably for 86 years while only being trained on greenhouse gas emissions data. We incorporate a fundamental physical law of advection-diffusion in PACER accounting for boundary conditions and empirically estimating the diffusion co-efficient and flow velocities from emissions data. PACER has been trained on 15 climate models provided by ClimateSet outperforming baselines across most of the climate models and advancing a new state of the art in a climate diagnostic task.
@@ 1 Introduction
To address this gap, we propose PACER, which treats climate emulation as a diagnostic-type prediction and integrate emissions data directly into the model’s training framework, to predict climate variables from a given parallel time-series of climate forcer emission maps (GHG and aerosols) allowing for more accurate simulation of future climate states under varying concentration scenarios.
Furthermore, we focus on two key phenomenon observed by our climate system i.e. advection and diffusion. In climate modeling, the advection-diffusion equation is fundamental for simulating the transport and dispersion of climate variables, such as temperature and moisture (Choi et al., 2023). PACER proves to be compute-efficient by modelling key physical law which reduces its dependence on large datasets making it data-efficient 1. Additionally, with generalizability inherent in the advection-diffusion equation, PACER generalizes across 15 climate models emulating surface temperature and precipitation stably for 86 years solely from emissions data (see Figure 1). Our contributions are as follows:
@@
    We propose PACER, a Neural ODE based climate emulator which models the advection-diffusion phenomenon by dynamically estimating the diffusion coefficient and flow velocities based on the input greenhouse gas concentrations.
@@
    Finally, we perform extensive experiments to show the generalization capabilities of PACER for emulating 15 climate models for 86 years at one time.
@@
Table 1: Computational Efficiency of PACER vs several Climate Emulators
@@ 3.3 Convolution Block Attention Module (CBAM)
Figure 2: Complete architectural pipeline of PACER. The model initializes the velocity fields and diffusion co-efficient from the input data of four gases i.e. $C0_{2},CH_{4},BC,SO_{2}$. After solving the advection-diffusion equation, channel and spatial attention extracts important spatial features to generate output maps of Temperature and Precipitation
@@ 4.1 Task
The goal of PACER is to emulate surface air temperature (TAS) and precipitation (PR) from climate forcer emission maps (CO2, CH4, SO2, BC) for a parallel time series of 2015-2100. We simulate the output of each climate model as single and super emulator, and also validate the generalisation of our methodology using zero-shot learning. We compare PACER against all baselines provided by ClimateSet under the same hyperparameter settings. We also compare against ACE (Watt-Meyer et al., 2023) and LUCIE Guan et al. (2024), two recent climate emulators. Since, they both are developed for different emulation task, we adopt their base architecture SFNO Bonev et al. (2023) and train it for the same task as ours. The details for adaptation of SFNO are given in Appendix A.2.
@@ 4.2 Dataset
We train PACER on a total of 15 climate models provided by ClimateSet Kaltenborn et al. (2023). ClimateSet compiles climate data from the Coupled Model Intercomparison Project Phase 6 (CMIP6) (Eyring et al., 2016) , incorporating climate model outputs from ScenarioMIP (O’Neill et al., 2016) and future emission trajectories of climate forcing agents from Input Datasets for Model Intercomparison Projects (Input4MIPs) (Durack et al., 2017). Each climate model has been standardized to a spatial resolution of 250km i.e. 96 $\times$ 144 grid points (latitude $\times$ longitude) with a monthly temporal resolution. Both input and output datasets consist of 86-year time-series data spanning four SSP scenarios (SSP1-2.6, SSP2-4.5, SSP3-7.0, SSP5-8.5) from 2015 to 2100. We use three scenarios namely SSP1-2.6, SSP3-7.0, SSP5-8.5 for training with a validation split of 0.1 and SSP2-4.5 for testing.
@@ 4.3 Evaluation Metrics
We evaluate PACER and all benchmarks using latitude-weighted Root Mean Square Error (RMSE) given in equation 19.
@@ 4.4 Single Emulator
For single emulator experiments, we trained all models for 25 epochs. We report RMSE for UNet, ConvLSTM, ClimaX, ClimaX_Frozen and SFNO. The training hyperparameters are all kept similar to those used in the original paper. PACER outperforms all models for emulating temperature across 13 climate models whereas ClimaX performs better on EC-Earth3 and TaiESM1 with lowest RMSE. For precipitation emulation, all models perform slightly worse with SFNO having lowest RMSE for simulating BCC-CSM2-MR and TaiESM1, ClimaX for AWI-CM-1-1-MR and CAS-ESM2-0, UNet performs best for INM-CM4-8 while PACER performs best for simulating remaining 10 climate models. Figure 3 shows the overall distribution of how each ML model performs across all 15 climate models in emulating temperature and precipitation. Detailed RMSE results for all 15 climate models for temperature and precipitation emulation are given in Appendix B in Figures 6 and 7.
Figure 3: RMSE distribution for climate variable predictions ccross different models. Each box represents the interquartile range (IQR) of RMSE values, capturing the spread and variability in prediction accuracy for key climate variables Temperature and Precipitation. The plot shows the consistent performance of PACER across all 15 climate models.
@@ 4.5 Super Emulator
Here, the term super emulator is used to train a single ML model on all of 15 climate models. This leads to the rich feature learning resulting in better generalization capabilities across different climate models. We use the same multihead decoder proposed in ClimateSet to train all ML models including PACER. We train all ML models for 100 epochs to keep the training regime computationally efficient with 2 convolutional layers and and 32 units decoder head. For super-emulator experiments we use a batch size of 1 for all models due to computational constraints.
For super emulation PACER outperforms all ML models on 13 climate models while ConvLSTM performs best for emulating EC-Earth3-Veg-LR and TaiESM1. The authors of ClimateSet Kaltenborn et al. (2023) suggest that during super emulation, smaller models demonstrate superior learning efficiency compared to larger models. This is because smaller models converge faster, allowing the model to learn patterns and relationships in the data more rapidly. We believe PACER being physically consistent and compute-efficient is able to learn complex climate features and outperform computationally intensive climate emulators.
@@
Climate Models PACER UNet ConvLSTM SFNO ClimaX ClimaXfrozen
@@
4.6 Generalization Capabilities of PACER as a Single Emulator
We perform generalization experiments to test the generalization capabilities of ML models on three climate models: AWI-CM-1-1-MR, MPI-ESM1-2-HR and FGOALS-f3-L. The RMSE results are shown in Tables 3, 4 and 5 for TAS (surface air temperature) and PR (precipitation) pre-trained on different climate models and tested on these three climate models. The metric for best performing model is emboldened. The pretrained climate model column shows which dataset the model was initially trained on before being tested on the either of the three climate models. Overall we observe that PACER, SFNO and ClimaX performs well for generalizing over different climate models with PACER outperforming on majority of the models. Here ClimaX does gain benefit from being pre-trained on a range of climate models initially. In future, we will test ClimaX generalization without using the pre-trained checkpoint to see how it compares against the smaller models such as unet and convlstm.
@@ 4.6 Generalization Capabilities of PACER as a Single Emulator
Pre-Trained Climate Model PACER UNet ConvLSTM SFNO ClimaX ClimaXfrozen TAS PR TAS PR TAS PR TAS PR TAS PR TAS PR AWI-CM-1-1-MR 0.184 0.499 0.289 0.571 0.451 0.622 0.200 0.501 0.207 0.498 0.412 0.707 BCC-CSM2-MR 0.247 0.620 0.275 0.653 0.466 0.696 0.230 0.599 0.232 0.617 0.432 0.753 CAS-ESM2-0 0.275 0.641 0.273 0.694 0.477 0.714 0.263 0.656 0.250 0.654 0.461 0.763 MRI-ESM2-0 0.205 0.620 0.276 0.656 0.456 0.697 0.218 0.635 0.223 0.643 0.410 0.756 NorESM2-MM 0.231 0.582 0.301 0.562 0.459 0.674 0.245 0.591 0.286 0.551 0.441 0.754
@@
Pre-Trained Climate Model PACER UNet ConvLSTM SFNO ClimaX ClimaXfrozen TAS PR TAS PR TAS PR TAS PR TAS PR TAS PR MPI-ESM1-2-HR 0.193 0.477 0.234 0.410 0.449 0.636 0.198 0.586 0.214 0.509 0.411 0.716 CNRM-CM6-1-HR 0.221 0.580 0.268 0.657 0.482 0.714 0.234 0.601 0.225 0.599 0.453 0.756 EC-Earth3 0.228 0.544 0.262 0.558 0.465 0.667 0.259 0.551 0.220 0.567 0.443 0.744 EC-Earth3-Veg-LR 0.241 0.554 0.270 0.560 0.471 0.660 0.224 0.515 0.233 0.551 0.456 0.744 TaiESM1 0.208 0.602 0.294 0.711 0.461 0.692 0.248 0.645 0.268 0.682 0.427 0.757
@@
Pre-Trained Climate Model PACER UNet ConvLSTM SFNO ClimaX ClimaXfrozen TAS PR TAS PR TAS PR TAS PR TAS PR TAS PR FGOALS-f3-L 0.184 0.559 0.241 0.562 0.485 0.652 0.253 0.561 0.218 0.573 0.456 0.729 GFDL-ESM4 0.207 0.563 0.321 0.716 0.484 0.697 0.271 0.600 0.330 0.708 0.468 0.762 INM-CM4-8 0.232 0.722 0.296 0.776 0.491 0.744 0.259 0.737 0.245 0.745 0.468 0.790 INM-CM5-0 0.212 0.574 0.277 0.756 0.488 0.739 0.261 0.730 0.250 0.725 0.459 0.785 MPI-ESM1-2-HR 0.209 0.696 0.283 0.701 0.482 0.714 0.257 0.701 0.250 0.691 0.445 0.760
@@
We utilize FDM for spatial discretization in PACER which divides the the physical space (in this case Earth atmosphere) into a grid of discrete points. Each grid point represents a specific location, and the value of the physical quantity (e.g., temperature) is computed at each point. The gridding at a lower resolution does induce additional errors. In future, we aim to test FVM and FEM to test if they results in smoother outputs and reduce errors.
@@ 6 Ablation Studies
To understand the importance of each component of PACER, we perform ablation studies across four climate models namely AWI-CM-1-1-MR, TaiESM1, EC-Earth3 and NorESM2-MM.
Advection-Only: For this study, we remove the empirical estimated diffusion term from PACER and only model the advection process using Neural ODE. The resulting RMSE for surface air temperature and precipitation increases deteriorating the model’s overall performance. The results show that missing approximation of diffusion has a greater effect on temperature as compared to precipitation, therefore determining that diffusion is critical in accurately simulating the transport of physical quantities like heat, moisture, and momentum.
@@ 7 Conclusion and Future Work
In this work, we present PACER, a physics and uncertainty aware climate emulator which accounts for Earth’s atmospheric advection-diffusion phenomenon. We incorporate a key physical law in PACER by solving a time-dependent partial different equation (PDE) using Neural ODE. Additionally, we encode periodic boundary conditions to avoid artificial edge effects that arise from rigid boundaries.
However, there are a few potential limitations of PACER which will be addressed as a possible future work. Our training regime only includes climate model data which is based on simulations and not considered entirely accurate. As a future work, we aim to extend training mechanism of PACER on both ERA5 weather and ClimateSet’s extensive climate data to enhance emulation accuracy. Additionally, PACER is trained on coarse resolution data which does not fully account for extreme events at regional level. Further, PACER is still limited in its ability to emulate precipitation stably for 86 years. These limitations can be addressed by training on high resolution data and encoding physical constraints such as energy, mass and water conservation in a loss function.
@@ A.1 Hyperparameters
Table 6: Training Hyperparameters for PACER
@@ A.3 Hardware and Software Requirements
We use PyTorch Paszke et al. (2019), Pytorch Lightning Falcon (2019), torchdiffeq Chen et al. (2018) for implementation of PACER. We train PACER on a single RTXA5000 with 24GB RAM. We perform all super emulator training experiments on a single NVIDIA DGX A100 with 80 GB RAM.
@@ Appendix B Results
Figure 6: RMSE results for Surface Air Temperature (TAS) Emulation for the projection of SSP2-4.5 (2015 – 2100). PACER outperforms all other ML models on 13 out of 15 Climate models namely:
Figure 7: RMSE results for Precipitation (PR) Emulation for the projection of SSP2-4.5 (2015 – 2100). PACER outperforms all other ML models on 9 out of 15 Climate models namely: